# CricaVPR: Cross-image Correlation-aware Representation Learning for Visual Place Recognition

### v1

@@ Abstract
Over the past decade, most methods in visual place recognition (VPR) have used neural networks to produce feature representations. These networks typically produce a global representation of a place image using only this image itself and neglect the cross-image variations (e.g. viewpoint and illumination), which limits their robustness in challenging scenes. In this paper, we propose a robust global representation method with cross-image correlation awareness for VPR, named CricaVPR. Our method uses the self-attention mechanism to correlate multiple images within a batch. These images can be taken in the same place with different conditions or viewpoints, or even captured from different places. Therefore, our method can utilize the cross-image variations as a cue to guide the representation learning, which ensures more robust features are produced. To further facilitate the robustness, we propose a multi-scale convolution-enhanced adaptation method to adapt pre-trained visual foundation models to the VPR task, which introduces the multi-scale local information to further enhance the cross-image correlation-aware representation. Experimental results show that our method outperforms state-of-the-art methods by a large margin with significantly less training time. Our method achieves 94.5% R@1 on Pitts30k using 512-dim global features. The code is released at https://github.com/Lu-Feng/CricaVPR.
@@ 1 Introduction
Figure 1: The Recall@1 and descriptors dimensionality comparison of different SOTA methods on Pitts30k. The GCL, NetVLAD, SFRS, and CricaVPR (Ours) all use PCA for dimensionality reduction. Our method achieves significantly higher Recall@1 than other methods with 512-dim compact global features.
@@
VPR is typically addressed as an image retrieval problem [12]. The place images are represented using global features and the similarity search is implemented in this feature space to return the matched place image. The global features are usually derived through the aggregation or pooling of local features, employing methods such as NetVLAD [5] or GeM [52] pooling. Such compact features are suitable for large-scale VPR. However, they lack robustness in challenging environments and are often susceptible to perceptual aliasing. An approach to improve robustness is to perform re-ranking by matching local features [26, 61], which incurs huge overhead in both runtime and memory footprint, i.e., not suitable for large-scale VPR. One problem that has been neglected is that existing methods produce the feature of an image only using this image itself (without cross-image interaction), which does not explicitly consider cross-image variations. To address this, our method attempts to use the cross-image variations as a cue to guide the representation learning and harvest useful information from other images when producing the feature of an image, making the output feature condition-invariant, viewpoint-invariant, and capable of addressing perceptual aliasing.
@@
Moreover, the recent visual foundation models [53, 71, 48] have achieved powerful performance. However, due to the particularity of the VPR task, directly using the pre-trained foundation model will encounter some problems. For example, the image features produced using pre-trained models tend to ignore some discriminative backgrounds, and are susceptible to interference from dynamic foregrounds (see Fig. 6 in experiments). Fine-tuning the model on VPR datasets can address this but tends to hurt the previously learned ability, i.e., catastrophic forgetting [19]. A promising way is to exploit parameter-efficient transfer learning (PETL) [28, 29]. However, the discriminative landmarks that need attention in VPR often occupy local regions of uncertain size in images, and most existing PETL methods use language-oriented adaptation modules to adapt the transformer model and lack the image-related (multi-scale) local priors for visual tasks (especially for VPR). This raises the need to develop a new adaptation method to introduce multi-scale local priors to the foundation model for VPR. Taking multi-scale variations into account is also an important complement to our cross-image correlation awareness.
@@
In this paper, we propose a novel method to learn Cross-image correlation-aware representation for VPR, abbreviated as CricaVPR. Our method first uses a backbone with the pooling module to yield initial feature representations. Then we use the cross-image encoder equipped with self-attention to calculate the correlation between multiple image representations within a batch to get final features. The images in a batch can be taken in the same place under different conditions (e.g. lighting) or from different viewpoints, or even captured from different places. This process allows each feature to enhance itself with useful information from others, thus producing condition-robust, viewpoint-robust, and discriminative representations. Meanwhile, we use the visual foundation model as the backbone in our architecture, and design a multi-scale convolution-enhanced adaptation method, in which we freeze the pre-trained foundation models and insert a few trainable lightweight adapters, to introduce the multi-scale local prior knowledge and adapt the foundation model for the VPR task.
@@
Our work brings the following contributions: 1) We propose a cross-image correlation-aware representation method, which uses self-attention to model the correlation between multiple image representations within a batch and make each feature more robust. Our method can also be combined with common aggregation methods to significantly boost performance. 2) We design a parameter-efficient adaptation method to adapt pre-trained models for VPR, in which the proposed multi-scale convolution adapter is used to introduce multi-scale local information to boost performance. 3) Extensive experiments on several VPR benchmark datasets show that our method can outperform the state-of-the-art (SOTA) methods by a large margin (see Fig. 1 for results on Pitts30k) with less training time.
@@ 2 Related Work
Two typical ways to alleviate this issue are to impose temporal consistency constraints and spatial consistency constraints. The former performs image sequence matching (i.e. utilize temporal continuity) [46, 17, 24, 41, 20] to realize robust VPR in challenging environments. The latter is often developed as a two-stage VPR system [25, 26, 7, 61, 72, 42, 43], which searches for top-k candidate images over the database using global features, then performs spatial consistency matching using local features to re-rank candidates. Different from these methods bringing additional constraints, runtime, and memory overhead, our model learns highly robust global representation via cross-image correlation awareness for global-retrieval-based VPR.
@@ 3 Methodology
Our method involves the Vision Transformer (ViT) model and the self-attention mechanism. So we first briefly review them in this section. Then we propose the cross-image correlation-aware representation method to describe place images. Finally, we present the multi-scale convolution-enhanced adaptation method to adapt the foundation model for VPR and the training strategy for fine-tuning.
@@ 3.1 Preliminary
The ViT model [18] and its variants have been applied for many computer vision tasks including VPR mainly due to its superior performance in modeling long-range dependencies. To process an input image with ViT, the image is initially divided into $N$ non-overlapping patches, which are then linearly projected into D-dim patch embeddings $x_{p}\in\mathcal{R}^{N\times D}$. Meanwhile, a learnable [class] token is prepended to $x_{p}$ to form $x_{0}=[x_{class};x_{p}]\in\mathcal{R}^{(N+1)\times D}$. To preserve the original positional information of each patch token, the corresponding positional embeddings are added to $x_{0}$ to get $z_{0}$, which is fed into a series of transformer encoder layers to yield the feature representation. A standard transformer encoder layer consists of three main components: the multi-head attention (MHA) layer, the MLP layer, and the LayerNormalization (LN) layer. The forward process of input $z_{l-1}$ passing through a transformer encoder layer to yield the output $z_{l}$ can be formulated as
@@
The MLP layer is made up of two fully connected layers, which are mainly used for feature nonlinearization and dimension conversion. Here we briefly overview the process of calculating the correlation and attention in the MHA layer. The input $X\in\mathcal{R}^{N\times D}$ is first linearly transformed to produce the queries $Q$, keys $K$, and values $V$. Then the attention among $Q,K$ and $V$ is computed using the Scaled Dot-Product Attention [60], denoted as
$\mathit{Attn}(Q,K,V)=\mathit{Softmax}\left(QK^{\top}/\sqrt{d}\right)V.$ (2)
The MHA first splits the queries, keys, and values into $h$ parts and performs attention for each part in parallel. The outputs of $h$ attention heads are then concatenated. The $N$ elements in the input sequence $X$ are correlated in this process. Specifically, we regard each element as a query, and the remaining elements as keys and values. For each query, we calculate the attention score between it and each key, establishing correlations between every query and all keys. These scores are then multiplied with their corresponding values, modeling dependencies among the $N$ elements.
There are two ways to yield global representations of places using the output of ViT. The first is to directly use the output class token as a global feature. The second is to reshape the output patch tokens as a feature map (similar to the output of CNN) to restore spatial position, and use the aggregation/pooling method (e.g. GeM [52]) to process it as a global feature. Both class token and GeM pooling are used to produce the place representation in our work.
@@ 3.2 Cross-image Correlation-aware Place Representation
Figure 3: The pipeline to produce the proposed cross-image correlation-aware representation. The cross-image encoder is the core component for modeling correlations between different image features in a batch. Note that we are correlating the $i$-th regional features of all images in a batch, not all regional features of an image.
@@
We design the pipeline to produce desired global features as shown in Fig. 3. The output patch tokens of a batch of images from the ViT backbone are reshaped as the $B\times W\times H\times D$-dim (i.e., batchbatch sizesize × weightweight × heightheight × tokentoken dimensiondimension) feature maps. We first use the spatial pyramid [36] to produce initial feature representations. The feature maps are split at three levels ($\times1$, $\times2$, and $\times3$). Then we use GeM pooling to process local (patch) features within the divided regions and get a total of 14 regional features of each image. Since the first level is a global aggregation, we directly use the class token to replace the GeM feature in this level for better performance. Next is the most critical step. We treat the $i$-th regional features of all images in a batch as a sequence of embedding vectors $f_{i}$, that is
@@
and feed the 14 sequences of embedding vectors into the cross-image encoder to model the correlation between the $i$-th regional features of all images in a batch. Instead of directly using the self-attention layer, the cross-image encoder is structured using two transformer encoder layers that also include the subsequent MLP layer and the skip connection for stable training and better performance. That is, we apply Eq. 1 to process each $f_{i}$ to correlate image representations in a batch. The 14 output regional features of each image are sequentially concatenated (i.e., flattened) as the final global representation of the image.
It should be noted that the retrieval process of our method is the same as the common global-retrieval-based method. That is, it uses the global feature of a single image for retrieval, which does not bring additional restrictions. Besides, we choose the spatial pyramid to produce the initial feature in order to subsequently correlate images at different scales, and the final sequential concatenation of regional features also introduces spatial position information to the global representations. In fact, our method can also use other aggregation methods to yield initial features, and boost the performance of these methods.
@@ 3.3 Multi-scale Convolution-enhanced Adaptation
Different from the vanilla adapter, our MulConv adapter adds a MulConv module between the (ReLU) activation layer and the up-projection layer. Meanwhile, there is a skip connection in parallel to the MulConv module. The MulConv module consists of three parallel convolutional paths of different scales ($\times1$, $\times3$, $\times5$). The $\times1$ convolution is also used before the $\times3$ and $\times5$ convolutions to reduce channel dimension. This design and the bottleneck structure of the adapter make our MulConv adapter still lightweight. The outputs of the three convolutional paths are concatenated to form the output of the MulConv module. Finally, the MulConv adapter is added in parallel to the MLP layer in each transformer block (i.e. transformer encoder layer) of the ViT backbone to achieve multi-scale convolution-enhanced adaptation, which can introduce proper (multi-scale) local priors to the model and improve performance for VPR. So the computation of each adapted transformer block can be denoted as
@@
Figure 4: Illustration of our multi-scale convolution-enhanced adaptation. (a) is a standard transformer block. (b) is the MulConv adapter. We add the MulConv adapter in parallel to the MLP layer in each transformer block to achieve our adaptation as in (c).
@@ 3.4 Training Strategy
Different from the common VPR works training models with weak supervision, we train our model on the GSV-Cities [1] dataset with full supervision. This dataset contains 560k images captured at 67k places with highly accurate labels. We follow the standard framework of this dataset and use the multi-similarity (MS) loss [63] with online hard mining strategy for training, as in the GSV-Cities [1] and MixVPR [2] works. The MS loss is computed as
$\begin{split}\mathcal{L}_{MS}=\frac{1}{m}\sum_{q=1}^{m}\bigg\{\frac{1}{\alpha}{\log\big[1+\sum_{p\in\mathcal{P}_{q}}e^{-\alpha(S_{qp}-\lambda)}}\big]\\ +\frac{1}{\beta}{\log\big[1+\sum_{n\in\mathcal{N}_{q}}e^{\beta(S_{qn}-\lambda)}\big]}\bigg\},\end{split}$ (5)
@@ 4.1 Datasets and Performance Evaluation
The experiments are conducted on several VPR benchmark datasets. These datasets exhibit viewpoint changes, condition changes, and the perceptual aliasing issue. Table 1 summarizes the key information of them. Pitts30k [57] contains a total of 30k database images in train, validation, and test sets. We use the test set to evaluate models. MSLS [64] consists of more than 1.6 million images captured in urban, suburban, and natural scenes from 30 cities on six continents over 7 years. We evaluate models on both MSLS-val and MSLS-challenge sets. The latter is an online test set without publicly released labels. Tokyo24/7 [58] contains about 76k reference images and 315 query images collected in urban scenes, and shows severe illumination (day/night) changes. We also use three challenging datasets: Nordland (with seasonal changes) [10], SVOX (cross-domain dataset) [11], and AmsterTime (with very long-term changes) [69]. More details are in Supplementary (Suppl.) Material.
The Recall@N (R@N) metric is used in our experiments to evaluate recognition performance. It is the percentage of queries for which at least one of the N retrieved database images is taken within a threshold of ground truth. We set the threshold to 25 meters and 40 for MSLS, 25 meters for Pitts30k, Tokyo24/7, and SVOX, $\pm 10$ frames for Nordland, unique counterpart for AmsterTime, following common evaluation procedures [64, 57, 58].

### v2

@@ Abstract
Over the past decade, most methods in visual place recognition (VPR) have used neural networks to produce feature representations. These networks typically produce a global representation of a place image using only this image itself and neglect the cross-image variations (e.g. viewpoint and illumination), which limits their robustness in challenging scenes. In this paper, we propose a robust global representation method with cross-image correlation awareness for VPR, named CricaVPR. Our method uses the attention mechanism to correlate multiple images within a batch. These images can be taken in the same place with different conditions or viewpoints, or even captured from different places. Therefore, our method can utilize the cross-image variations as a cue to guide the representation learning, which ensures more robust features are produced. To further facilitate the robustness, we propose a multi-scale convolution-enhanced adaptation method to adapt pre-trained visual foundation models to the VPR task, which introduces the multi-scale local information to further enhance the cross-image correlation-aware representation. Experimental results show that our method outperforms state-of-the-art methods by a large margin with significantly less training time. The code is released at https://github.com/Lu-Feng/CricaVPR.
@@ 1 Introduction
Figure 1: The Recall@1 and descriptors dimensionality comparison of different methods on Pitts30k. The GCL, NetVLAD, SFRS, and CricaVPR (Ours) all use PCA for dimensionality reduction. Our method can achieve significantly higher Recall@1 than other methods with 512-dim compact global features.
@@
VPR is typically addressed as an image retrieval problem [12]. The place images are represented using global features and the similarity search is implemented in this feature space to return the matched place image. The global features are usually derived through the aggregation (pooling) of local features, employing methods such as NetVLAD [5] or GeM [53] pooling. Such compact features are suitable for large-scale VPR. However, they lack robustness in challenging environments and are often susceptible to perceptual aliasing. A way to improve robustness is to perform re-ranking by matching local features [26, 62], which incurs huge overhead in runtime and memory footprint, making it difficult to achieve large-scale VPR. One problem that has been neglected is that existing methods produce the feature of an image only using this image itself (without cross-image interaction), which does not explicitly consider cross-image variations. To address this, our method attempts to use the cross-image variations as a cue to guide the representation learning and harvest useful information from other images when producing the feature of an image, making the output feature condition-invariant, viewpoint-invariant, and capable of addressing perceptual aliasing.
@@
Moreover, the recent visual foundation models [54, 72, 49] have achieved powerful performance. However, due to the particularity of the VPR task, directly using the pre-trained foundation model will encounter some problems. For example, the image features produced using pre-trained models tend to ignore some discriminative backgrounds, and are susceptible to interference from dynamic foregrounds (see Fig. 6 in experiments). Fine-tuning the model on VPR datasets can address this but tends to hurt the previously learned ability, i.e., catastrophic forgetting [19]. A promising way is to exploit parameter-efficient transfer learning (PETL) [28, 29]. However, the discriminative landmarks that need attention in VPR often occupy local regions of uncertain size in images, and most existing PETL methods use language-oriented adaptation modules to adapt the transformer model and lack the image-related (multi-scale) local priors for visual tasks (especially for VPR). This raises the need to develop a new adaptation method to introduce multi-scale local priors to the foundation model for VPR.
@@
In this paper, we propose a novel method to learn Cross-image correlation-aware representation for VPR, abbreviated as CricaVPR. Our method first uses a backbone with a pooling module to yield initial feature representations. Then we use a cross-image encoder equipped with the attention to calculate the correlation between multiple image representations within a batch to get final features. The images in a batch can be taken in the same place under different conditions (e.g. lighting) or from different viewpoints, or even captured from different places. This process allows each feature to enhance itself with useful information from others, thus producing condition-robust, viewpoint-robust, and discriminative representations. Meanwhile, we use the visual foundation model as the backbone in our architecture, and design a multi-scale convolution-enhanced adaptation method, in which we freeze the pre-trained foundation models and insert a few trainable lightweight adapters, to introduce the multi-scale local prior knowledge and adapt the foundation model for the VPR task.
@@
Our work brings the following contributions: 1) We propose a cross-image correlation-aware representation method, which uses the attention mechanism to model the correlation between multiple image representations within a batch and make each feature more robust. 2) We design a parameter-efficient adaptation method to adapt pre-trained models for VPR, in which the proposed multi-scale convolution adapter is used to introduce multi-scale local information to boost performance. 3) Extensive experiments on the benchmark datasets show that our method can outperform the state-of-the-art (SOTA) methods by a large margin with less training time. The results on Pitts30k that best reflect the advantages of our method are shown in Fig. 1.
@@ 2 Related Work
Two typical ways to alleviate this issue are to impose temporal consistency constraints and spatial consistency constraints. The former performs image sequence matching (i.e. utilize temporal continuity) [47, 17, 24, 41, 20] to realize robust VPR in challenging environments. The latter is often developed as a two-stage VPR system [25, 26, 7, 42, 62, 73, 43, 44], which searches for top-k candidate images over the database using global features, then performs spatial consistency matching using local features to re-rank candidates. Different from these methods bringing additional constraints, runtime, and memory overhead, our model learns highly robust global representation via cross-image correlation awareness for global-retrieval-based VPR.
@@ 3 Methodology
Our method involves the Vision Transformer (ViT) and the attention mechanism used in it. So we first briefly review them in this section. Then we propose the cross-image correlation-aware representation method to describe place images. Finally, we present the multi-scale convolution-enhanced adaptation method to adapt the foundation model for VPR and the training strategy for fine-tuning.
@@ 3.1 Preliminary
The ViT model [18] and its variants have been applied for many computer vision tasks including VPR mainly due to its superior performance in modeling long-range dependencies. To process an input image with ViT, the image is initially divided into $N$ non-overlapping patches, which are then linearly projected into $D$-dim patch embeddings $x_{p}\in\mathcal{R}^{N\times D}$. Meanwhile, a learnable [class] token is prepended to $x_{p}$ to form $x_{0}=[x_{class};x_{p}]\in\mathcal{R}^{(N+1)\times D}$. To preserve the original positional information of each patch token, the corresponding positional embeddings are added to $x_{0}$ to get $z_{0}$, which is fed into a series of transformer encoder layers to yield the feature representation. A transformer encoder layer consists of three main components: the multi-head attention (MHA) layer, the MLP layer, and the LayerNormalization (LN) layer. The forward process of input $z_{l-1}$ passing through a transformer encoder layer to yield the output $z_{l}$ can be formulated as
@@
The MLP layer is made up of two fully connected layers, which are mainly used for feature nonlinearization and dimension conversion. Here we briefly overview the process of calculating the correlation and attention in the MHA layer. The input sequence is first linearly transformed to produce the queries $Q$, keys $K$, and values $V$. Then the attention among $Q,K$ and $V$ is computed using the Scaled Dot-Product Attention [61], denoted as
[…]
The MHA utilizes different learnable linear projections to generate the queries, keys, and values $h$ times and performs attention for each set of projections in parallel. Specifically, we first compute the attention scores between each query and all keys, establishing the correlations between them. These scores are then multiplied with the corresponding values to model dependencies among these tokens. Finally, the outputs of $h$ attention heads are concatenated (and once again projected). All tokens/elements in the input sequence are correlated in this process. In the next section, we will also use this attention mechanism to compute the across-image correlation.
There are two ways to yield global representations of places using the output of ViT. The first is to directly use the output class token as a global feature. The second is to reshape the output patch tokens as a feature map (similar to the output of CNN) to restore the spatial position, and use the aggregation/pooling method (e.g. GeM [53]) to process it as a global feature. Both the class token and GeM pooling are used to produce the place representation in our work.
@@ 3.2 Cross-image Correlation-aware Place Representation
Figure 3: The pipeline to produce the proposed cross-image correlation-aware representation. The cross-image encoder is the core component for modeling correlations between different image features in a batch. Note that we are correlating the $i$-th regional features of all images in a batch, not all regional features of an image. Besides, the cross-image encoder consists of 2 stacked vanilla transformer encoder layers [61] with the LN layer behind the MHA/MLP layer, which is different from that in ViT [18] (LN is before MHA/MLP).
@@
We design the pipeline to produce desired global features as shown in Fig. 3. The output patch tokens of a batch of images from the ViT backbone are reshaped as the $B\times W\times H\times D$-dim (i.e., $batch$ $size\times weight\times height\times token$ $dimension$) feature maps. We first use the spatial pyramid [36] to produce initial feature representations. The feature maps are split at three levels (1$\times$1, 2$\times$2, and 3$\times$3). Then we use GeM pooling to process local (patch) features within the divided regions and get a total of 14 regional features of each image. Since the first level is a global aggregation, we directly use the class token to replace the GeM feature in this level for better performance. Next is the most critical step. We treat the $i$-th regional features of all images in a batch as a sequence of embedding vectors $f_{i}$, that is
@@
and feed the 14 sequences of embedding vectors into a cross-image encoder to model the correlation between the $i$-th regional features of all images in a batch. That is, we apply the cross-image encoder to process each $f_{i}$ to correlate images in a batch. Instead of directly using the attention (MHA) layer, the cross-image encoder is structured using two (vanilla) transformer encoder layers [61] that also include the MLP layer, LN layer, and skip connection for stable training and better performance. The 14 output regional features of each image are sequentially concatenated (i.e., flattened) and L2-normalized as the final global representation of the image.
It should be noted that the retrieval process of our method is the same as the common global-retrieval-based method. That is, it uses the global feature of a single image for retrieval. Besides, we choose the spatial pyramid to produce the initial feature in order to subsequently correlate images at different scales, and the final sequential concatenation of regional features also introduces spatial position information to the global representations. In fact, our method can also use other aggregation methods to yield initial features, and boost the performance of these methods.
@@ 3.3 Multi-scale Convolution-enhanced Adaptation
Different from the vanilla adapter, our MulConv adapter adds a MulConv module between the (ReLU) activation layer and the up-projection layer. This module consists of three parallel convolutional paths of different scales (1$\times$1, 3$\times$3, 5$\times$5). The 1$\times$1 convolution is also used before the 3$\times$3 and 5$\times$5 convolutions to reduce channel dimension. This design and the bottleneck structure of the adapter make our MulConv adapter still lightweight. The outputs of the three convolutional paths are concatenated to form the output of the MulConv module. Besides, there is a skip connection in parallel to the MulConv module. Finally, the MulConv adapter is added in parallel to the MLP layer (multiplied by a scaling factor $s$) in each transformer block (i.e. transformer encoder layer) of the ViT backbone to achieve multi-scale convolution-enhanced adaptation, which can introduce proper (multi-scale) local priors to the model and improve performance for VPR. So the computation of each adapted transformer block can be denoted as
@@
Figure 4: Illustration of our multi-scale convolution-enhanced adaptation. (a) is a transformer block in ViT. (b) is the MulConv adapter. We add the MulConv adapter in parallel to the MLP layer in each transformer block to achieve our adaptation as in (c).
@@ 3.4 Training Strategy
We train our model on the GSV-Cities [1] dataset with full supervision. This dataset contains 560k images captured at 67k places with highly accurate labels. We follow the standard framework of this dataset and use the multi-similarity (MS) loss [64] with online hard mining strategy for training. The MS loss is computed as
$\begin{split}\mathcal{L}_{MS}=\frac{1}{B}\sum_{q=1}^{B}\bigg{\{}\frac{1}{\alpha}{\log\big{[}1+\sum_{p\in\mathcal{P}_{q}}e^{-\alpha(S_{qp}-\lambda)}}\big{]}\\ +\frac{1}{\beta}{\log\big{[}1+\sum_{n\in\mathcal{N}_{q}}e^{\beta(S_{qn}-\lambda)}\big{]}}\bigg{\}},\end{split}$ (5)
@@ 4.1 Datasets and Performance Evaluation
The experiments are conducted on several VPR benchmark datasets. These datasets exhibit viewpoint changes, condition changes, and the perceptual aliasing issue. Table 1 summarizes the key information of them. Pitts30k [58] mainly shows large viewpoint changes. MSLS [65] consists of images captured in urban, suburban, and natural scenes over 7 years, and covers various visual changes. Tokyo24/7 [59] exhibits severe illumination (day/night) changes. We also use three challenging datasets: Nordland (with seasonal changes) [10], SVOX (cross-domain dataset) [11], and AmsterTime (with very long-term changes) [70]. More details are in Supplementary (Suppl.) Material.
The Recall@N (R@N) metric is used in our experiments to evaluate recognition performance. It is the percentage of queries for which at least one of the N retrieved database images is taken within a threshold of ground truth. We set the threshold to 25 meters and 40 for MSLS, 25 meters for Pitts30k, Tokyo24/7, and SVOX, ±$\pm 10$ frames for Nordland, unique counterpart for AmsterTime, following common evaluation procedures [65, 58, 59].